OpenStudy (anonymous):

how many sides does a regular polygon have if each interior angle measures 144 degrees

OpenStudy (anonymous):

my say is 5

OpenStudy (anonymous):

5 or 6 if that helps

OpenStudy (anonymous):

decagon

OpenStudy (anonymous):

180(n-2) = 144n n = number of sides the formula is \[\frac{ 180*(n-2) }{ n }\] = intern angles

OpenStudy (anonymous):

\[\frac{ 180*(n-2) }{ n } = interior angles\]